# Netan Dogra

### v1

@@ 6.1. Galois cohomology of number fields
    $\dim H^{1}_{f}(G_{\mathbb{Q}_{p}},W)-\dim H^{1}_{f}(G_{\mathbb{Q},T},W)$
    $=$ $\dim H^{0}(G_{\mathbb{R}},W)-\dim H^{1}_{f}(G_{\mathbb{Q},T},W^{*}(1)).$
@@
    $\dim\mathrm{Ker}(H^{1}(G_{K,T},W)\to\oplus_{v\in T}H^{1}(G_{v},W))$
    $=$ $\dim\mathrm{Ker}(H^{2}(G_{K,T},W^{*}(1))\to\oplus_{v\in T}H^{2}(G_{v},W^{*}(1))).$
@@
$\dim H^{1}_{f}(G_{\mathbb{Q}},H^{m}_{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r))$
$=$ $\dim H^{1}_{f}(G_{\mathbb{Q}_{p}},H^{m}_{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r)))-\dim H^{0}(G_{\mathbb{R}},H^{m}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r)))$
$-\dim\mathrm{Hom}_{G_{\mathbb{Q}}}(H^{m}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r)),\mathbb{Q}_{p}(1)).$
@@ 6.2. The Chabauty–Coleman–Kim sets $X(\mathbb{Q}_{p})_{n}$
Let $X$ be a smooth projective geometrically irreducible curve over a number field $K$, of genus $g>1$. We assume we have a given $K$-rational points $b\in X(K)$. Fix a prime $v$ of $K$ lying above a prime $p$ of $\mathbb{Q}$. The Chabauty–Coleman–Kim method, as developed in [Kim09], [KT08] and [Kim05], determines subsets $X(K_{v})_{n}\subset X(K_{v})$, containing the set of rational points $X(K)$, which are defined in terms of the $\mathbb{Q}_{p}$-unipotent fundamental group of $X$, which we now define.
@@
$\rho_{n}:\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)\to U_{n}(b)(\mathbb{Q}_{p}),$
@@
$\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)\to U(\mathbb{Q}_{p}),$
@@
[…]
commute. By universal properties, $U_{n}(b)$ inherits an action of $\operatorname{Gal}(\overline{K}|K)$ from $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{K}},b)$ making the homomorphism $\rho_{n}$ Galois-equivariant. Hence we obtain morphisms
$j_{n,K}:X(K)\to H^{1}(G_{K},U_{n}(b)).$
@@
    For $G=G_{\mathbb{Q},S}$ or $G_{\mathbb{Q}_{v}}$ for a prime $v$ of $\mathbb{Q}$, $H^{1}(G,U_{n}(b))$ is isomorphic to the $\mathbb{Q}_{p}$-points of a scheme of finite type over $\mathbb{Q}_{p}$, in such a way that all the maps $\mathrm{loc}_{v}$ and $s_{a,b,m,n*}$ are induced from morphisms of schemes.
@@
    $H^{1}_{f}(G_{\mathbb{Q}_{v}},U_{n}(b))$ is (the $\mathbb{Q}_{p}$-points of) a $\mathbb{Q}_{p}$-subvariety of $H^{1}(G_{K_{v}},U_{n}(b))$, and moreover this $\mathbb{Q}_{p}$-subvariety is isomorphic to an affine space.
We define $\mathrm{Sel}(U_{n}(b))\subset H^{1}(G_{\mathbb{Q},S},U_{n}(b))$ to be the fibre product, over $H^{1}(G_{\mathbb{Q},S},U_{1}(b))\times\prod_{v\in S}H^{1}(G_{\mathbb{Q}_{v}},U_{n}(b))$, of $H^{1}(G_{\mathbb{Q},S},U_{n}(b))$ with the following subschemes:
$H^{1}_{f}(G_{\mathbb{Q}_{p}},U_{n}(b))\subset H^{1}(G_{\mathbb{Q}_{p}},U_{n}(b)),$
$j_{n,\mathbb{Q}_{v}}(X(\mathbb{Q}_{v}))\subset H^{1}(G_{\mathbb{Q}_{v}},U_{n}(b)),$ $v\in S-\{p\}$
$\mathrm{Im}(\mathrm{Jac}(X)(K)\otimes\mathbb{Q}_{p})\subset H^{1}(G_{K,S},U_{1}(b)).$
@@
$\operatorname{rk}J(\mathbb{Q})+\sum_{i=2}^{n}\dim H^{1}_{f}(G_{\mathbb{Q}},\mathrm{gr}_{i}U_{n})<\sum_{i=1}^{n}\dim H^{1}_{f}(G_{\mathbb{Q}_{p}},\mathrm{gr}_{i}U_{n}).$
@@
$\dim H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}V)<\frac{(3g-2)(g+1)}{2}-\operatorname{rk}J(\mathbb{Q}).$
@@ 6.3. Dimension bounds via descent
$\dim H^{1}(\mathcal{O}_{K,S},\wedge^{2}J[2])=\operatorname{rk}\mathcal{O}_{K_{f}^{(2)},S}^{\times}+\operatorname{rk}\mathcal{O}_{K,S}^{\times}-\operatorname{rk}\mathcal{O}_{K_{f},S}^{\times}.$
@@ 7. Computation of the generalised height
$\sum h_{v}(x)=h(\mathrm{AJ}(x-\infty),\mathrm{AJ}(x-\infty)).$
@@
In fact, there will be many such height functions $h_{v}$ that we want to use, and it is better to linearly combine them all to construct vector space-valued heights: specifically we will construct local heights with values in $H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$, so that the identity (16) is taking place in
$\oplus_{v\in S}H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)).$
The reader familiar with the theory of $p$-adic height pairings over number fields may think of $\oplus_{v\in S}H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ as a generalisation of the space of functionals on idele class characters.
We first recall the definition and basic properties, following [BD21, §3].
@@
$H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V/\mathbb{Q}_{p}(1))\to H^{1}(G_{\mathbb{Q}_{v}},U_{2})$
@@
$\widetilde{h}_{v}:H^{1}(G_{\mathbb{Q}_{v}},U_{2})\to H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V/\mathbb{Q}_{p}(1)),$
@@
$\widetilde{h}_{p}:H^{1}_{f}(G_{\mathbb{Q}_{p}},U_{2})\to H^{1}_{f}(G_{\mathbb{Q}_{p}},\wedge^{2}V/\mathbb{Q}_{p}(1))$
@@
in $J(\mathbb{Q})^{\otimes 2}\otimes\mathbb{Q}$, then there is an element $\xi\in H^{1}_{f,S_{0}}(G_{\mathbb{Q}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ such that
@@
$\mathbb{Q}_{p}[X(\mathbb{Q})]\to\oplus_{v}H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
@@
To get a symmetric map (i.e. a map factoring through $\mathrm{Sym}^{2}J(\mathbb{Q})$) we can remove the alternating part as follows. For $M_{1},M_{2}\in M(\mathbb{Q}_{p},V)$, define $M_{1}\wedge M_{2}\in M(\mathbb{Q}_{p},V^{\oplus 2},\wedge^{2}V/\mathbb{Q}_{p}(1))$ to be the quotient of $M_{1}\otimes M_{2}$ by $\mathrm{Sym}^{2}V_{p}J\oplus\mathbb{Q}_{p}(1)$. We define
$h_{v}(z_{1},z_{2}):=\widetilde{h}_{v}(\mathcal{E}_{2}(z_{1},z_{2}))-\frac{1}{2}\widetilde{h}_{v}(\mathcal{E}_{1}(z_{1},z_{2})\wedge\mathcal{E}_{1}(z_{2},w(z_{2}))).$
@@
$h:\mathbb{Q}_{p}[X(\mathbb{Q})\otimes X(\mathbb{Q})]\to\oplus_{v}H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
@@
Note that when $v\neq p$, $h_{v}=\widetilde{h}_{v}$ since $H^{1}(G_{\mathbb{Q}_{v}},V)=0$. The question of how to compute the local heights $\widetilde{h}_{v}$ for $v\neq p$ is a special case of the question of how to describe the action of monodromy on the unipotent fundamental group and associated path torsors. This was studied by Kim and Tamagawa [KT08] and also in recent work with Betts [BD19]. For studying the BMSST curve, we will only need the following result.
@@ 7.1. Computing $X(\mathbb{Q}_{2})_{2}$
denote the generalised height with respect to a basepoint $b$. For $v\neq 2$, let $\Omega_{v}\subset H^{1}(G_{\mathbb{Q}_{v}},\wedge^{2}T_{2}V/\mathbb{Q}_{p}(1))$ denote the image of $X(\mathbb{Q}_{v})$ under $j_{v}$. Suppose for simplicity that all the $\alpha_{v}$ are trivial (this is the setting we will be in in the case of the BMSST curve). Let $X$ be a hyperelliptic curve of genus $g$ and Mordell–Weil rank $g\leq r<g^{2}$. Suppose that we have verified the rank part of the Bloch–Kato conjecture for $H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}V_{p}J)$, so that $\dim H^{1}_{f}(G_{\mathbb{Q}},\wedge^{2}V_{p}J)=(g-1)(g+2)/2$ and $X(\mathbb{Q}_{p})_{2}$ is finite. For a divisor $D=\sum n_{i}z_{i}$, we will slightly bizarrely write $h_{v}(D)$ to mean $\sum_{i}n_{i}h_{v}(z_{i})$. Define $\Theta$ to be the map

### v2

@@ 6.1. Galois cohomology of number fields
    $\dim H^{1}_{f}(\mathbb{Q}_{p},W)-\dim H^{1}_{f}(G_{\mathbb{Q},T},W)$
    $=$ $\dim H^{0}(\mathbb{R},W)-\dim H^{1}_{f}(G_{\mathbb{Q},T},W^{*}(1)).$
@@
    $\dim\mathrm{Ker}(H^{1}(G_{K,T},W)\to\oplus_{v\in T}H^{1}(K_{v},W))$
    $=$ $\dim\mathrm{Ker}(H^{2}(G_{K,T},W^{*}(1))\to\oplus_{v\in T}H^{2}(K_{v},W^{*}(1))).$
@@
$\dim H^{1}_{f}(\mathbb{Q},H^{m}_{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r))$
$=$ $\dim H^{1}_{f}(\mathbb{Q}_{p},H^{m}_{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r)))-\dim H^{0}(\mathbb{R},H^{m}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r)))$
$-\dim H^{0}(\mathbb{Q},H^{m}(X_{\overline{\mathbb{Q}}},\mathbb{Q}_{p}(r-1))^{*}).$
@@ 6.2. The Chabauty–Coleman–Kim sets $X(\mathbb{Q}_{p})_{n}$
Let $X$ be a smooth projective geometrically irreducible curve over $\mathbb{Q}$, of genus $g>1$. Let $b\in X(\mathbb{Q})$. Fix a prime $p$. The Chabauty–Coleman–Kim method, as developed in [Kim09], [KT08] and [Kim05], determines subsets $X(\mathbb{Q}_{p})_{n}\subset X(\mathbb{Q}_{p})$, containing the set of rational points $X(\mathbb{Q})$, which are defined in terms of the $\mathbb{Q}_{p}$-unipotent fundamental group of $X$, which we now define.
@@
$\rho_{n}:\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},b)\to U_{n}(b)(\mathbb{Q}_{p}),$
@@
$\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},b)\to U(\mathbb{Q}_{p}),$
@@
${\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},b)}$${U_{n}(b)}$${U(\mathbb{Q}_{p})}$
commute. By universal properties, $U_{n}(b)$ inherits an action of $\operatorname{Gal}(\overline{\mathbb{Q}}|\mathbb{Q})$ from $\pi_{1}^{\mathrm{\acute{e}t}}(X_{\overline{\mathbb{Q}}},b)$ making the homomorphism $\rho_{n}$ Galois-equivariant. Hence we obtain morphisms
$j_{n,K}:X(\mathbb{Q})\to H^{1}(\mathbb{Q},U_{n}(b)).$
@@
    For $G=G_{\mathbb{Q},S}$ or $\operatorname{Gal}(\overline{\mathbb{Q}}_{v}|\mathbb{Q}_{v})$ for a prime $v$ of $\mathbb{Q}$, $H^{1}(G,U_{n}(b))$ is isomorphic to the $\mathbb{Q}_{p}$-points of a scheme of finite type over $\mathbb{Q}_{p}$, in such a way that all the maps $\mathrm{loc}_{v}$ and $s_{a,b,m,n*}$ are induced from morphisms of schemes.
@@
    $H^{1}_{f}(G_{\mathbb{Q}_{v}},U_{n}(b))$ is (the $\mathbb{Q}_{p}$-points of) a $\mathbb{Q}_{p}$-subvariety of $H^{1}(\mathbb{Q}_{v},U_{n}(b))$, and moreover this $\mathbb{Q}_{p}$-subvariety is isomorphic to an affine space.
We define $\mathrm{Sel}(U_{n}(b))\subset H^{1}(G_{\mathbb{Q},S},U_{n}(b))$ to be the fibre product, over $H^{1}(G_{\mathbb{Q},S},U_{1}(b))\times\prod_{v\in S}H^{1}(\mathbb{Q}_{v},U_{n}(b))$, of $H^{1}(G_{\mathbb{Q},S},U_{n}(b))$ with the following subschemes:
$H^{1}_{f}(\mathbb{Q}_{p},U_{n}(b))\subset H^{1}(\mathbb{Q}_{p},U_{n}(b)),$
$j_{n,\mathbb{Q}_{v}}(X(\mathbb{Q}_{v}))\subset H^{1}(\mathbb{Q}_{v},U_{n}(b)),$ $v\in S-\{p\}$
$\mathrm{Im}(\mathrm{Jac}(X)(\mathbb{Q})\otimes\mathbb{Q}_{p})\subset H^{1}(G_{\mathbb{Q},S},U_{1}(b)).$
@@
$\operatorname{rk}J(\mathbb{Q})+\sum_{i=2}^{n}\dim H^{1}_{f}(\mathbb{Q},\mathrm{gr}_{i}U_{n})<\sum_{i=1}^{n}\dim H^{1}_{f}(\mathbb{Q}_{p},\mathrm{gr}_{i}U_{n}).$
@@
$\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}V)<\frac{(3g-2)(g+1)}{2}-\operatorname{rk}J(\mathbb{Q}).$
@@ 6.3. Dimension bounds via descent
$\dim H^{1}(\mathcal{O}_{K,S},\wedge^{2}J[2])$ $=\operatorname{rk}\mathcal{O}_{K_{f}^{(2)},S}^{\times}+\operatorname{rk}\mathcal{O}_{K,S}^{\times}-\operatorname{rk}\mathcal{O}_{K_{f},S}^{\times}$
$+\dim\mathrm{Cl}(\mathcal{O}_{K_{f}^{(2)},S})[2]-\dim\mathrm{Cl}(\mathcal{O}_{K_{f},S})[2]+\dim\mathrm{Cl}(\mathcal{O}_{K,S})[2].$
@@ 7. Computation of the generalised height
$\sum h_{v}(x)=h(\mathrm{AJ}(x-\{\infty\}),\mathrm{AJ}(x-\{\infty\})).$
@@
In fact, there will be many such height functions $h_{v}$ that we want to use, and it is better to linearly combine them all to construct vector space-valued heights: specifically we will construct local heights with values in $H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$, so that the identity (16) is taking place in
$\oplus_{v\in S}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)).$
The reader familiar with the theory of $p$-adic height pairings over number fields may think of $\oplus_{v\in S}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ as a generalisation of the space of functionals on idele class characters.
We first recall the definition and basic properties. We refer the reader to [BD21, §3] for more details.
@@
$H^{1}(\mathbb{Q}_{v},\wedge^{2}V/\mathbb{Q}_{p}(1))\to H^{1}(\mathbb{Q}_{v},U_{2})$
@@
$\widetilde{h}_{v}:H^{1}(\mathbb{Q}_{v},U_{2})\to H^{1}(\mathbb{Q}_{v},\wedge^{2}V/\mathbb{Q}_{p}(1)),$
@@
$\widetilde{h}_{p}:H^{1}_{f}(\mathbb{Q}_{p},U_{2})\to H^{1}_{f}(\mathbb{Q}_{p},\wedge^{2}V/\mathbb{Q}_{p}(1))$
@@
in $J(\mathbb{Q})^{\otimes 2}\otimes\mathbb{Q}$, then there is an element $\xi\in H^{1}_{f,S_{0}}(\mathbb{Q},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ such that
@@
$\mathbb{Q}_{p}[X(\mathbb{Q})]\to\oplus_{v}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
@@
To get a symmetric map (i.e. a map factoring through $\mathrm{Sym}^{2}J(\mathbb{Q})$) we can remove the alternating part as follows. First recall (from [BD21]) that the generalised height is really a map on the set $M(\mathbb{Q}_{p},V_{p}J,\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$ of equivalence classes of mixed extensions of $\mathbb{Q}_{p}$-Galois representations with graded pieces $\mathbb{Q}_{p},V_{p}J$ and $\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)$. The map on $X(K)$ (for $K=\mathbb{Q}$ or $\mathbb{Q}_{v}$) is then defined by sending $x$ to the generalised height evaluated at the mixed extension $E(b,x)_{\mathbb{Q}_{2}}$, where $E(b,x)_{\mathbb{Q}_{2}}$ is defined as follows. Let $E_{n}(b,x)$ be the Galois representation defined in section 5. Let $E_{n}(b,x)_{\mathbb{Q}_{2}}:=E_{n}(b,x)\otimes_{\mathbb{Z}_{2}}\mathbb{Q}_{2}$. Then $E_{2}(b,x)_{\mathbb{Q}_{2}}$ is an extension of $E_{1}(b,x)_{\mathbb{Q}_{2}}$ by
$\mathrm{Ker}(H^{1}_{\mathrm{\acute{e}t}}(X_{\overline{K}},\mathbb{Q}_{p})^{\otimes 2}\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}H^{2}_{\mathrm{\acute{e}t}}(X_{\overline{K}},\mathbb{Q}_{p}))^{*}.$
Define $E(b,x)$ to be $E_{2}(b,x)_{\mathbb{Q}_{2}}/\mathrm{Sym}^{2}V_{p}(J)$.
Using the notion of equivariant generalised heights (see [BD21, §4]) one can also evaluate such a generalised height on a mixed extension with graded pieces $\mathbb{Q}_{p},V_{p}J^{\oplus n}$ and $\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1)$. Now for $M_{1},M_{2}\in M(\mathbb{Q}_{p},V)$, define $M_{1}\wedge M_{2}\in M(\mathbb{Q}_{p},V^{\oplus 2},\wedge^{2}V/\mathbb{Q}_{p}(1))$ to be the quotient of $M_{1}\otimes M_{2}$ by $\mathrm{Sym}^{2}V_{p}J\oplus\mathbb{Q}_{p}(1)$. We define
$h_{v}(z_{1},z_{2}):=\widetilde{h}_{v}(E(z_{1},z_{2}))-\frac{1}{2}\widetilde{h}_{v}(E_{1}(z_{1},z_{2})_{\mathbb{Q}_{2}}\wedge E_{1}(z_{2},w(z_{2}))_{\mathbb{Q}_{2}}).$
@@
$h:\mathbb{Q}_{p}[X(\mathbb{Q})\times X(\mathbb{Q})]\to\oplus_{v}H^{1}(\mathbb{Q}_{v},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))/H^{1}(G_{\mathbb{Q},S},\wedge^{2}V_{p}J/\mathbb{Q}_{p}(1))$
@@
Note that when $v\neq p$, $h_{v}=\widetilde{h}_{v}$ since $H^{1}(\mathbb{Q}_{v},V)=0$. The question of how to compute the local heights $\widetilde{h}_{v}$ for $v\neq p$ is a special case of the question of how to describe the action of monodromy on the unipotent fundamental group and associated path torsors. This was studied by Kim and Tamagawa [KT08] and also in recent work with Betts [BD19]. For studying the BMSST curve, we will only need the following result.
@@ 7.1. Computing $X(\mathbb{Q}_{2})_{2}$
denote the generalised height with respect to a basepoint $b$. For $v\neq 2$, let $\Omega_{v}\subset H^{1}(\mathbb{Q}_{v},\wedge^{2}T_{2}V/\mathbb{Q}_{p}(1))$ denote the image of $X(\mathbb{Q}_{v})$ under $j_{v}$. Suppose for simplicity that all the $\alpha_{v}$ are trivial (this is the setting we will be in in the case of the BMSST curve). Let $X$ be a hyperelliptic curve of genus $g$ and Mordell–Weil rank $g\leq r<g^{2}$. Suppose that we have verified the rank part of the Bloch–Kato conjecture for $H^{1}_{f}(\mathbb{Q},\wedge^{2}V_{p}J)$, so that $\dim H^{1}_{f}(\mathbb{Q},\wedge^{2}V_{p}J)=(g-1)(g+2)/2$ and $X(\mathbb{Q}_{p})_{2}$ is finite. For a divisor $D=\sum n_{i}z_{i}$, we will slightly bizarrely write $h_{v}(D)$ to mean $\sum_{i}n_{i}h_{v}(z_{i})$. Define $\Theta$ to be the map